### v1

# Finitary conditions in special semidirect products

Abstract.
We initiate the study of finitary conditions for monoids arising as expansions via the semidirect product construction of a monoid $M$ acting on the left of its power set (which we term ‘special semidirect products’, and denote by $\mathcal{S}(M)$). Specifically, we examine the condition of weak left coherence (every finitely generated left ideal has a finite presentation as a left act); the related conditions of property (L), left ideal Howson, finitely left equated; and the left-right dual properties of weak right coherence, property (R), right ideal Howson and finitely right equated. Each of these conditions is preserved under retract, from which it is immediate that if $\mathcal{S}(M)$ satisfies one of our finitary conditions, then so must $M$, but the converse is not true. Indeed, we show that $\mathcal{S}(M)$ satisfies property (L) (or property (R)) if and only if $M$ is finite. We provide a characterisation of the monoids $M$ such that $\mathcal{S}(M)$ is left (respectively, right) ideal Howson, and give sufficient conditions for $\mathcal{S}(M)$ to be finitely left (respectively, right) equated and hence weakly left (respectively, right) coherent.
Key words and phrases:
Expansions, special semidirect products, finitary conditions
2020 Mathematics Subject Classification:
20M10, 20M30
1. Introduction
This paper initiates the study of finitary conditions (that is, conditions satisfied by every finite monoid) in the context of certain expansions of monoids. Monoid expansions provide systematic ways to build a new monoid from a given input monoid, often by inflating the set of idempotents in some structured manner. They were introduced in [1] to facilitate the application of techniques in finite semigroup theory to infinite monoids and semigroups. Most well-studied examples of expansions have the property that when restricting the input to an interesting class of monoids (e.g. groups, cancellative monoids, unipotent monoids), the output of the expansion is a monoid with desirable structural properties. For example, the Szendrei expansion [18] maps: groups to inverse monoids [1], right cancellative monoids to left ample monoids [5] and unipotent monoids to (what are now called) left restriction monoids [6]. In fact, [1] used an earlier formulation, known as the prefix expansion, which is equivalent to the Szendrei expansion in the case the monoid is a group; [14] investigates the prefix expansion of an arbitrary monoid. The underlying construction used in these expansions is a semidirect product formed from the monoid $M$ acting on a particular semilattice of subsets of $M$. It is then natural to consider the ‘parent’ expansion where the semilattice contains all subsets of $M$, and this is our focus here. Specifically, this paper considers the expansion formed via the action of a monoid $M$ on its power set $\mathcal{P}(M)$ viewed as a monoid (indeed, a semilattice) with respect to the operation of union, and with identity element $\emptyset$, with left action of $M$ on $\mathcal{P}(M)$ given by $(m,X)\mapsto mX=\{mx:x\in X\}$ for all $X\in\mathcal{P}(M)$ and all $m\in M$. Throughout the paper we will write $\mathcal{S}(M)$ to denote $\mathcal{P}(M)\rtimes M$ with product given by $(X,x)(Y,y)=(X\cup xY,xy)$, and with identity element $(\emptyset,1)$, where $1$ is the identity element of $M$, and we shall refer to monoids of this form as ‘special semidirect products’.
Our interest in special semidirect products comes from several directions. First, $\mathcal{S}(M)$ contains the Szendrei expansion of $M$ as a monoid subsemigroup, and (as with the Szendrei expansion) if $G$ is a group then $\mathcal{S}(G)$ is an inverse monoid. If $M=F_{X}$ is the free group on a set $X$, we remark that the monoid $\mathcal{S}(F_{X})$ contains: the free inverse, free ample and free left ample monoids on the same set $X$. For a monoid $M$ which is not a group, the semidirect product $\mathcal{S}(M)$ is no longer inverse (indeed, it is not even regular; see Lemma 3.3 below for details). It is then natural to ask under which circumstances $\mathcal{S}(M)$ belongs to one of several more general classes of monoids (e.g. left/right restriction, ample, adequate, abundant).
Second, we note that any inverse monoid satisfies several finitary conditions, including weak left (respectively, right) coherence [2]. A monoid $M$ is said to be weakly left (respectively, right) coherent if every finitely generated left (respectively, right) ideal of $M$ is finitely presented as a left (respectively, right) $M$-act. A classical result of Wheeler [19], interpreted for acts in [9], says that a monoid is left coherent if and only if the class of existentially closed left acts is axiomatisable. Inspired by this, weak left coherency for monoids, and the related finitary conditions we consider here, have also been formulated in terms of the model-theoretic notion of axiomatisability: see [8, 17, 4]. In previous work [13] the construction $\mathcal{S}(M)$ has been utilised to exhibit some interesting examples of monoids that are neither left nor right coherent. We have remarked that the monoid $\mathcal{S}(F_{X})$ contains the free inverse, free ample and free left ample monoids on set $X$, and it is known that each of the latter is weakly left and right coherent [2, 3, 11]; cf. the table in [13]. One is then drawn to ask (see Question 8.6 of [13]): for which monoids $M$ is $\mathcal{S}(M)$ weakly left (respectively right) coherent? It is known that $M$ is weakly left coherent if and only if it is simultaneously left ideal Howson (the intersection of any two finitely generated left ideals is finitely generated) and finitely left equated (the left annihilator congruences of $M$ are finitely generated as left congruences). We therefore examine the behaviour of these two conditions and the closely related conditions of (R) and (L) with respect to the expansion $\mathcal{S}$.
The paper is structured as follows. In Section 2 we outline the preliminary definitions and results required for our paper and in Section 3 we prove some foundational results on special semidirect products. Our main results are in Section 4. We give precise characterisations of the monoids $M$ for which (R), (L) (Subsection 4.1) or the left and right ideal Howson properties hold in $\mathcal{S}(M)$ (Subsection 4.2). We exhibit some necessary conditions for $\mathcal{S}(M)$ to be finitely right or left equated and provide some configurations of elements in $M$ that prevent $\mathcal{S}(M)$ from being finitely right or left equated (Subsection 4.3). In particular, we apply our results to demonstrate that if $M$ is a right (left) LCM monoid then $\mathcal{S}(M)$ is weakly right (left) coherent. We provide a number of examples and counterexamples to illustrate and separate the properties under consideration. We finish with some open questions.
2. Preliminaries
In this section we give the necessary specialist background to follow this paper. We refer the reader to [15] for a fuller account of the semigroup theoretic notions we discuss. Throughout the paper $M$ denotes a monoid with identity element $1$.
2.1. Right congruences
An equivalence relation $\rho$ on $M$ is a right congruence if for all $a,b,c\in M$ if $a\,\rho\,b$ then $ac\,\rho\,bc$. Every right ideal of $M$ determines a Rees right congruence, but not all right congruences correspond to right ideals. If $W\subseteq M\times M$ then we denote the smallest right congruence containing $W$ by $\rho_{W}$ and refer to this as the right congruence generated by $W$. The following result is standard and may be found in [16].
Proposition 2.1.
Let $W\subseteq M\times M$ and let $a,b\in M$. Then $a\,\rho_{W}\,b$ if and only if there is a sequence
$a=c_{1}t_{1},\,d_{1}t_{1}=c_{2}t_{2},\,\cdots,\,d_{n}t_{n}=b,$
where $n\geq 0$, $(c_{i},d_{i})\in W\cup W^{-1}$ and $t_{i}\in M$, for $1\leq i\leq n$.
A sequence as in Proposition 2.1 is called a $W$-sequence of length $n$; we interpret $n=0$ as saying that $a=b$. The dual result holds for left congruences; we denote the left congruence generated by $W$ by $\lambda_{W}$.
2.2. Classes of monoids determined by extensions of Green’s relation $\mathcal{L}$
We recall several well-known classes of semigroups that may be viewed as generalisations of regular (every $\mathcal{L}$-class and every $\mathcal{R}$-class contains an idempotent) and inverse (regular and the idempotents form a semilattice) monoids determined by relations that extend Green’s relations $\mathcal{L}$ and $\mathcal{R}$ with respect to a distinguished semilattice of idempotents. We recap the relevant definitions extending $\mathcal{L}$; for brevity we omit the natural left-right dual definitions extending $\mathcal{R}$.
Definition 2.2.
(Right annihilator congruence, $\mathcal{L}^{*}$, $\tilde{\mathcal{L}}_{E}$, right abundant, right Fountain) For $a\in M$ let
$\mathbf{r}(a):=\{(u,v)\in M\times M:au\,=\,av\}.$
It is straightforward to check that $\mathbf{r}(a)$ is a right congruence; we call this the right annihilator congruence of $a$. We denote by $\mathcal{L}^{*}$ the equivalence relation on $M$ defined (for all $a,b\in M$) by $a\,\mathcal{L}^{*}\,b$ if $\mathbf{r}(a)=\mathbf{r}(b)$. The relation $\mathcal{L}^{*}$ is a generalisation of Green’s $\mathcal{L}$-relation in that it contains $\mathcal{L}$ and is a right congruence. Let $E(M)$ denote the set of idempotents of $M$, and let $E\subseteq E(M)$. The pre-order $\leq_{\widetilde{\mathcal{L}}_{E}}$ on $M$ is defined by the rule that for all $a,b\in M$:
$a\leq_{\widetilde{\mathcal{L}}_{E}}b\mbox{ if and only if }\{e\in E:be=b\}\subseteq\{e\in E:ae=a\}.$
The associated equivalence relation is denoted by $\widetilde{\mathcal{L}}_{E}$. Thus, $a\,\widetilde{\mathcal{L}}_{E}\,b$ if and only if $a$ and $b$ have the same right identities from $E$. The relation $\widetilde{\mathcal{L}}_{E}$ is a generalisation of both Green’s $\mathcal{L}$-relation and the relation $\mathcal{L}^{*}$: indeed, we have that $\mathcal{L}\subseteq\mathcal{L}^{*}\subseteq\mathcal{L}_{E}$, and if $M$ is regular we have $\mathcal{L}=\widetilde{\mathcal{L}}_{E(M)}$. In general, however, unlike the relations $\mathcal{L}$ and $\mathcal{L}^{*}$, the relation $\widetilde{\mathcal{L}}_{E}$ need not be a right congruence. We say that $M$ is right abundant (respectively, right Fountain) if every $\mathcal{L}^{*}$–class (respectively, $\widetilde{\mathcal{L}}$–class) of $M$ contains an idempotent.
Definition 2.3.
(Right $E$-Ehresmann, $E$-adequate, $E$-ample and $E$-restriction monoids) Suppose now that $E\subseteq E(M)$ is a semilattice (for example, one could take $E=\{1\}$). Then each $\widetilde{\mathcal{L}}_{E}$-class of $M$ contains at most one idempotent of $E$, and the monoid $M$ is said to be right $E$-Ehresmann if every $\widetilde{\mathcal{L}}_{E}$-class of $M$ contains a (unique) element of $E$ and $\widetilde{\mathcal{L}}_{E}$ is a right congruence. In a right $E$-Ehresmann monoid then, there is a natural unary operation determined by $x\mapsto x^{*}$ for all $x\in M$ where $x^{*}$ denotes the unique element of $E$ in the $\widetilde{\mathcal{L}}_{E}$ class of $x$. If $M$ is right $E$-Ehresmann and $\widetilde{\mathcal{L}}_{E}=\mathcal{L}^{*}$, then $M$ is said to be right $E$-adequate. We say that $M$ is right $E$-ample (respectively, right $E$-restriction) if it is right $E$-adequate (respectively right $E$-Ehresmann) and satisfies the right ample identity $ea=a(ea)^{*}$ for all $a\in M$ and all $e\in E$. We denote by $\sigma_{E}$ the least (two-sided) congruence on $M$ identifying all elements of $E$ or simply $\sigma$ in the case where $E=E(M)$. An inverse (respectively, right $E$-ample; right $E$-restriction) monoid $M$ is said to be proper if $\mathcal{L}\,\cap\,\sigma$ (respectively, $\mathcal{L}^{*}\,\cap\sigma_{E}$; $\tilde{\mathcal{L}}_{E}\,\cap\sigma_{E}$) is trivial.
Remark 2.4.
[…]
Remark 2.5.
If $E(M)$ is itself a semilattice, then we may take $E=E(M)$ in each of the above definitions; in this case we suppress the $E$ denoting the corresponding relation by $\tilde{\mathcal{L}}$ and using the terminology right Ehresmann, right adequate, right ample and right restriction. For consistency we keep this terminology although we note that if $M$ is right $E$-ample then it turns out that we must in fact have $E=E(M)$. To see this, let $e\in E(M)$ and notice that since $e\,\widetilde{\mathcal{L}}_{E}\,e^{*}$ we have $ee^{*}=e$ giving $ee=ee^{*}$ so that (using that $e\,\mathcal{L}^{*}_{E}\,e^{*}$) $e^{*}e=e^{*}e^{*}=e^{*}$; applying the right ample identity then gives $e^{*}=e^{*}e=e(e^{*}e)^{*}=e(e^{*})^{*}=ee^{*}=e$ and hence $e\in E$.
2.3. A suite of finitary conditions
Weak right coherence is a finitary condition for monoids (in the sense that every finite monoid is weakly coherent) and is defined as follows:
Definition 2.6.
(Weak right coherence) A monoid $M$ is weakly right coherent if every finitely generated right ideal of M is finitely presented as a right $M$-act.
In practice, we shall make use of the following characterisation of weak right coherence.
Theorem 2.7.
[10, Corollary 3.3] A monoid $M$ is weakly right coherent if and only if
  1. (1)
    $aM\cap bM$ is finitely generated;
  2. (2)
    $\mathbf{r}(a)$ is a finitely generated right congruence on $M$.
It will be convenient to have some terminology corresponding to the conditions of the previous theorem.
Definition 2.8.
(Right ideal Howson and finitely right equated monoids) A monoid with the property that the intersection of any two principal (or equivalently two finitely generated) right ideals is finitely generated is said to be right ideal Howson. A monoid $M$ with the property that each right annihilator congruence of the form $\mathbf{r}(a)$ where $a\in M$ is finitely generated are said to be finitely right equated.
Another related finitary condition arises as follows:
Definition 2.9.
(Condition (R)) For $a,b\in M$ let
$\mathbf{R}(a,b):=\{(u,v)\in M\times M:au\,=\,bv\}\subseteq M\times M.$
It is easy to see that $\mathbf{R}(a,b)$ is a right subact of $M\times M$. We say that $M$ satisfies condition (R) if for all $a,b\in M$ the right subact $\mathbf{R}(a,b)$ is finitely generated.
Each of these properties behaves well with respect to retraction:
Theorem 2.10.
[4, Corollary 3.5, Corollary 4.12, Theorem 5.5], [11, Theorem 6.2] and [12, Theorem 2.3] Let $M$ be a monoid and $S$ be a retract of $M$. If $M$ is right ideal Howson (respectively, finitely right equated; weakly right coherent, satisfies condition (R)), then so is $S$.
We note that the term ‘finitely right aligned’ used elsewhere in the literature coincides with the property of being right ideal Howson for monoids. (Right ideal Howson semigroups need not be finitely aligned, however [3].)
For right abundant monoids, the notions of weak right coherence and right ideal Howson coincide:
Proposition 2.11.
[10, Lemma 3.5] Every right abundant monoid is finitely right equated.
3. Special semidirect products
We recall briefly that an expansion of monoids is a functor $\mathcal{F}$ from the category of monoids to the category of monoids such that there is a natural transformation $\rho$ from $\mathcal{F}$ to the identity functor, with surjective components, that is: for any two monoids $M,N$ there exist surjective morphisms $\rho_{M}:\mathcal{F}(M)\rightarrow M$ and $\rho_{N}:\mathcal{F}(N)\rightarrow N$ such that for any morphism $\varphi:M\rightarrow N$, the following diagram commutes:
[…]
As mentioned in the introduction $\mathcal{S}:M\rightarrow\mathcal{S}(M)$ is an expansion of monoids, as is easily seen by considering the surjection $(A,a)\mapsto a$ onto the second co-ordinate. We now record some basic facts about the structure of $\mathcal{S}(M):=\mathcal{P}(M)\rtimes M$
The following notational conventions will be useful.
Notation 3.1.
For $a\in M$ and $A\subseteq M$ we write $a^{-1}A$ and $\hat{A}$ to denote the sets
$a^{-1}A$ $:=$ $\{u\in M:au\in A\}$
$\hat{A}$ $:=$ $\{e\in E(M):eA\subseteq A\}.$
Remark 3.2.
For $a\in M$ and $X,Y\subseteq M$, $aY\subseteq X$ if and only if $Y\subseteq a^{-1}X$. In particular for any $X\subseteq M$ we have:
(3.1) $X\subseteq a^{-1}(aX)$ $=$ $\{u:u\in M,au\in aX\},$
(3.2) $a(a^{-1}X)$ $=$ $\{au:u\in M,au\in X\}=aM\cap X\subseteq X.$
  1. (1)
    It follows from (3.1) that $a^{-1}(aX)=X$ for all $a\in M$ and all $X\subseteq M$ if and only if $M$ is left cancellative. (In general the inclusion $\{b\}\subseteq a^{-1}(\{ab\})$ can be strict.)
  2. (2)
    It follows from (3.2) that $a(a^{-1}X)=X$ for all $a\in M$ and all $X\subseteq M$ if and only if $M$ is a group. (In general the inclusion $a(a^{-1}M)=aM\subseteq M$ can be strict.)
Lemma 3.3.
Let $M$ be a monoid.
  1. (1)
    The set of idempotents of $\mathcal{S}(M)$ is $\{(A,e):e\in\hat{A}\}$ and this is a semilattice precisely if $1$ is the unique idempotent of $M$. In general, $\mathcal{E}:=\{(A,1):A\subseteq M\}$ is a semilattice of idempotents.
  2. (2)
    The monoid $\mathcal{S}(M)$ is regular if and only if $M$ is a group, in which case $\mathcal{S}(M)$ is a proper inverse monoid (and hence, in particular, weakly left and right coherent).
Proof.
(1) The idempotents of $\mathcal{S}(M)$ are easily seen to be the elements of the form $(A,e)$ where $e$ is an idempotent of $M$ and $A\subseteq M$ satisfies $eA\subseteq A$. In particular $(\emptyset,1)$ is the identity element of $\mathcal{S}(M)$, where $1$ is the identity element of $M$. If $(A,e)$ and $(B,f)$ are idempotents of $\mathcal{S}(M)$, then $(A,e)(B,f)=(A\cup eB,ef)$ and $(B,f)(A,e)=(B\cup fA,fe)$ need not be idempotent, and need not be equal. If $1$ is the only idempotent of $M$, then $\mathcal{E}$ is the set of all idempotents of $\mathcal{S}(M)$ and, in any case, this is easily seen to be a semilattice since $(A,1)(B,1)=(A\cup B,1)=(B\cup A,1)=(B,1)(A,1)$. Conversely, suppose that the set of idempotents of $\mathcal{S}(M)$ is a semilattice. For any $e,f\in E(M)$ note that $(\emptyset,e)$ and $(\emptyset,f)$ are idempotents, and since these commute we must have $ef=fe$. Thus $E(M)$ must be a semilattice. Next, $(\{e\},ef)=(\{e\},e)(\emptyset,f)=(\emptyset,f)(\{e\},e)=(\{fe\},fe)$. Dually, $(\{f\},fe)=(\{ef\},ef)$ so that since $fe=ef$ we obtain that $e=f$.
(2) Suppose $\mathcal{S}(M)$ is regular and let $x\in M$. By assumption there exists $(Y,y)\in\mathcal{S}(M)$ with
$(\{1\},x)=(\{1\},x)(Y,y)(\{1\},x)=(\{1,xy\}\cup xY,xyx).$
Thus $xy=1$ and as this holds for every $x\in M$ we deduce that $M$ is a group. Conversely, if $M$ is a group, then it is easy to see that $\mathcal{S}(M)$ has a particularly nice structure; it is an inverse monoid with $(X,x)^{-1}=(x^{-1}X,x^{-1})$. Further, since $\mathcal{S}(M)$ is a semidirect product of a semilattice by a group it follows from classical results that it is proper [15]. It is instructive to see how this transpires: for any $(A,a),(B,b)\in\mathcal{S}(M)$ we have that $(A,a)\,\sigma\,(B,b)$ if and only if $a=b$, $(A,a)\,\mathcal{R}\,(B,b)$ if and only if $A=B$ and $(A,a)\,\mathcal{L}\,(B,b)$ if and only if $a^{-1}A=b^{-1}B$. Finally, since $\mathcal{S}(M)$ is inverse, it is weakly coherent by Proposition 2.11. ∎
We next consider the situations in which $\mathcal{S}(M)$ lies in the class of left abundant, left $\mathcal{E}$-Ehresmann, left $\mathcal{E}$-restriction or left $\mathcal{E}$-ample monoids.
Proposition 3.4.
Let $M$ be a monoid, $\mathcal{E}=\{(F,1):F\subseteq M\}$ and $(A,a),(B,b)\in\mathcal{S}(M)$.
  1. (1)
    $(A,a)\,\mathcal{R}^{*}\,(B,b)$ if and only if $A=B$ and $a\,\mathcal{R}^{*}\,b$.
  2. (2)
    $(A,a)\,\widetilde{\mathcal{R}}\,(B,b)$ if and only if $A=B$ and $a\,\widetilde{\mathcal{R}}_{\hat{A}}\,b$.
  3. (3)
    $(A,a)\,\widetilde{\mathcal{R}}_{\mathcal{E}}\,(B,b)$ if and only if $A=B$.
  4. (4)
    $\mathcal{R}^{*}=\widetilde{\mathcal{R}}_{\mathcal{E}}$ in $\mathcal{S}(M)$ if and only if $M$ is right cancellative.
  5. (5)
    $\mathcal{S}(M)$ is left abundant if and only if for each $A\subseteq M$ and $a\in M$ there exists an idempotent $e\in\hat{A}$ with $a\,\mathcal{R}^{*}\,e$.
  6. (6)
    $\mathcal{S}(M)$ is always a proper left $\mathcal{E}$-restriction monoid and so certainly left $\mathcal{E}$-Ehresmann; it is left $\mathcal{E}$-ample if and only if $M$ is right cancellative.
Proof.
(1) Suppose first that $(A,a)\,\mathcal{R}^{*}\,(B,b)$. Since $(A,1)(A,a)=(\emptyset,1)(A,a)$ and $(B,1)(B,b)=(\emptyset,1)(B,b)$ it follows that $A=A\cup B=B$. If $ua=va$ then $(\emptyset,u)(A,a)=(\emptyset,v)(A,a)$ giving $(\emptyset,u)(B,b)=(\emptyset,v)(B,b)$, and hence $ub=vb$. A dual argument shows that if $ub=vb$ then $ua=va$. Thus $a\,\mathcal{R}^{*}\,b$. Conversely, if $A=B$ and $a\,\mathcal{R}^{*}\,b$ then it is easy to see that $(U,u)(A,a)=(V,v)(A,a)$, if and only if $(U,u)(B,b)=(V,v)(B,b)$.
[…]
(3) Restricting attention to the idempotents of $\mathcal{E}$, it is clear that $(E,1)$ is a left identity for $(A,a)$ if and only if $E\subseteq A$, and hence the set of idempotents of $\mathcal{E}$ acting identically on the left of $(A,a)$ will be equal to the corresponding set of idempotents for $(B,b)$ if and only if $A=B$.
(4) Let $a\in M$ and consider $(\{1\},a)\in\mathcal{S}(M)$. It is easy to see that the idempotent left identities of $(\{1\},a)$ are $(\emptyset,1)$ and $(\{1\},1)$. Thus for all $a\in M$ we have $(\{1\},a)\,\widetilde{\mathcal{R}}_{\mathcal{E}}\,(\{1\},1)$. If $\mathcal{R}^{*}=\widetilde{\mathcal{R}}_{\mathcal{E}}$ it then follows from part (1) that $a\,\mathcal{R}^{*}\,1$ for all $a\in M$. Thus $ua=va$ if and only if $u=v$, that is, $M$ is right cancellative. Conversely, if $M$ is right cancellative, then $E(M)=\{1\}$ so that $\mathcal{E}=E(\mathcal{S}(M))$ and $a\,\mathcal{R}^{*}\,b$ holds for all $a,b\in M$. From parts (1) and (3) we have $\mathcal{R}^{*}=\widetilde{\mathcal{R}}$$\widetilde{\mathcal{R}}_{\mathcal{E}}$.
(5) By part (1) the $\mathcal{R}^{*}$-class of $(A,a)$ contains an idempotent $(E,e)$ if and only if $A=E$, $e\in\hat{A}$ and $a\,\mathcal{R}^{*}\,e$.
[…]
$(A,a)(B,b)^{+}$ $=$ $(A,a)(B,1)=(A\cup aB,a)=(A\cup aB,1)(A,a)$
$=$ $(A\cup aB,ab)^{+}(A,a)=((A,a)(B,b))^{+}(A,a).$
Thus $\mathcal{S}(M)$ is left $\mathcal{E}$-restriction. Recall that $\sigma_{\mathcal{E}}$ is the least congruence identifying all idempotents of $\mathcal{E}$. Since all idempotents in $\mathcal{E}$ have second component $1$, it is easy to see that if $(A,a)\,\sigma_{\mathcal{E}}\,(B,b)$ then we must have $a=b$. From part (3) we also have that $(A,a)\,\widetilde{\mathcal{R}}_{\mathcal{E}}\,(B,b)$ if and only if $A=B$. Thus $\widetilde{\mathcal{R}}_{\mathcal{E}}\cap\sigma_{\mathcal{E}}$ is trivial and hence $\mathcal{S}(M)$ is proper left $\mathcal{E}$-restriction. By part (4) we have seen that $\widetilde{\mathcal{R}}_{\mathcal{E}}$$\mathcal{R}^{*}$ if and only if $M$ is right cancellative; thus $\mathcal{S}(M)$ is proper left $\mathcal{E}$-ample if and only if $M$ is right cancellative. ∎
By Proposition 3.4, left abundance of $M$ is a necessary condition for $\mathcal{S}(M)$ to be left abundant. Likewise, by the left-right dual to Theorem 2.10, $M$ being finitely left equated is a necessary condition for $\mathcal{S}(M)$ to be finitely left equated. Neither condition is sufficient however, as we shall see in Example 4.21.
We now consider the situations in which $\mathcal{S}(M)$ lies in the class of right abundant, right $\mathcal{E}$-Ehresmann, right $\mathcal{E}$-restriction or right $\mathcal{E}$-ample monoids.
Proposition 3.5.
Let $M$ be a monoid, $\mathcal{E}=\{(E,1):E\subseteq M\}$ and $(A,a),(B,b)\in\mathcal{S}(M)$.
  1. (1)
    $(A,a)\,\widetilde{\mathcal{L}}\,(B,b)$ if and only if $a^{-1}A=b^{-1}B$ and $a\,\widetilde{\mathcal{L}}\,b$.
  2. (2)
    $(A,a)\,\mathcal{L}^{*}\,(B,b)$ if and only if $a^{-1}A=b^{-1}B$ and $a\,\mathcal{L}^{*}\,b$.
  3. (3)
    $(A,a)\,\widetilde{\mathcal{L}}_{\mathcal{E}}\,(B,b)$ if and only if $a^{-1}A=b^{-1}B$.
  4. (4)
    $\mathcal{L}^{*}=\widetilde{\mathcal{L}}_{\mathcal{E}}$ in $\mathcal{S}(M)$ if and only if $M$ is left cancellative.
  5. (5)
    $\mathcal{S}(M)$ is right abundant if and only if $M$ is right abundant.
  6. (6)
    Each $\tilde{\mathcal{L}}_{\mathcal{E}}$–class of $\mathcal{S}(M)$ contains a unique idempotent from $\mathcal{E}$, and $\mathcal{S}(M)$ is a right $\mathcal{E}$-Ehresmann monoid (or right $\mathcal{E}$-adequate) if and only if $M$ is left cancellative, in which case $\mathcal{E}=E(\mathcal{S}(M))$.
  7. (7)
    $\mathcal{S}(M)$ is right ample if and only if $M$ is a group (in which case $\mathcal{S}(M)$ is a proper inverse monoid).
Proof.
(1) Suppose first that $(A,a)\,\widetilde{\mathcal{L}}\,(B,b)$. Since $(A,a)(a^{-1}A,1)=(A,a)$ we must have $(B,b)(a^{-1}A,1)=(B,b)$, which implies that $b(a^{-1}A)\subseteq B$, which in turn implies that $a^{-1}A\subseteq b^{-1}B$. Thus with the dual argument we have that $a^{-1}A=b^{-1}B$. Next, suppose that $ae=a$. Since $(A,a)(\emptyset,e)=(A,e)$ we also require that $(B,b)(\emptyset,e)=(B,b)$, that is $be=b$, and it follows that $a\,\widetilde{\mathcal{L}}\,b$.
Conversely, suppose that $a^{-1}A=b^{-1}B$ and $a\,\widetilde{\mathcal{L}}\,b$, and let $(E,e)$ be an idempotent with $(A,a)(E,e)=(A,a)$. Since $A\cup aE=A$, we have that $aE\subseteq A$, or equivalently $E\subseteq a^{-1}A=b^{-1}B$, giving $bE\subseteq B$. Further, since $ae=a$ and $a\,\widetilde{\mathcal{L}}\,b$, we have $be=b$. Hence $(B,b)(E,e)=(B,be)=(B,b)$ and it follows that $(A,a)\,\widetilde{\mathcal{L}}\,(B,b)$.
(2) Suppose first that $(A,a)\,\mathcal{L}^{*}\,(B,b)$. Then $(A,a)\,\widetilde{\mathcal{L}}\,(B,b)$ so that by (1) we have $a^{-1}A=b^{-1}B$. Next, suppose that $ax=ay$. Since $(A,a)(\emptyset,x)=(A,a)(\emptyset,y)$ we also require that $(B,b)(\emptyset,x)=(B,b)(\emptyset,y)$, that is $bx=by$. Thus with the dual argument we obtain $a\,\mathcal{L}^{*}\,b$.
Conversely, suppose that $a^{-1}A=b^{-1}B$ and $a\,\mathcal{L}^{*}\,b$. Let $(A,a)(X,x)=(A,a)(Y,y)$. Then $A\cup aX=A\cup aY$ and $ax=ay$. From the latter we immediately deduce that $bx=by$. Meanwhile, the former gives $aX\setminus A=aY\setminus A$. Let $p\in B\cup bX$; we show that $p\in B\cup bY$. Clearly this is true if $p\in B$, so we assume $p\notin B$. Then $p=bu$ for some $u\in X$. If $au\in A$ then $u\in a^{-1}A=b^{-1}B$ so that $p\in B$, a contradiction. Thus $au\in aX\setminus A=aY\setminus A$, giving $au=av$ and then $p=bu=bv$ for some $v\in Y$, since $a\,\mathcal{L}^{*}\,b$. Thus in this case also $p\in B\cup bY$. Together with the dual argument we obtain $B\cup bX=B\cup bY$ from which it follows that $(B,b)(X,x)=(B,b)(Y,y)$.
[…]
(4) It is immediate from (2) and (3) that $\mathcal{L}^{*}=\mathcal{L}_{\mathcal{E}}$ if and only if all elements of $M$ are $\mathcal{L}^{*}$-related or, in other words, if and only if $M$ is left cancellative.
(5) It is immediate from part (2) that $(A,a)\,\mathcal{L}^{*}\,(E,e)$ if and only if $a\,\mathcal{L}^{*}\,e$ and $e^{-1}E=a^{-1}A$. Since $(E,e)$ being idempotent implies that $e$ is idempotent, $\mathcal{S}(M)$ right abundant implies that $M$ is right abundant. Conversely, suppose that $M$ is right abundant. Then for all $a\in M$ there exists an idempotent $e\in M$ such that for all $u,v\in M$, $au=av$ if and only if $eu=ev$. Since $e^{2}=e$ note that $e$ must be an idempotent right identity for $a$. Now taking $E=a^{-1}A$ it is straightforward to check that $a^{-1}A=e^{-1}E$ (since $u\in a^{-1}A$ if and only if $au=aeu\in A$ if and only if $eu\in a^{-1}A=E$ if and only if $u\in e^{-1}E$) and $(E,e)$ is idempotent (since $eE=e(a^{-1}A)=e((ae)^{-1}A)=e(e^{-1}(a^{-1}A))=e(e^{-1}E)\subseteq E$). It then follows from part (2) that $(A,a)\,\mathcal{L}^{*}\,(E,e)$.
[…]
$(at)^{-1}(A\cup aT)=t^{-1}(a^{-1}A\cup a^{-1}(aT))=t^{-1}(a^{-1}A\cup T)=t^{-1}(b^{-1}B\cup b^{-1}(bT))=(bt)^{-1}(B\cup bT),$
[…]
(7) From the previous part, we may define a unary operation $(A,a)^{*}=(a^{-1}A,1)$, and if $M$ is left cancellative then $\mathcal{S}(M)$ is right $\mathcal{E}$-Ehresmann with respect to this operation. However, $\mathcal{S}(M)$ need not be right $\mathcal{E}$-restriction. Indeed $(A,1)(B,b)=(A\cup B,b)$ but
$\begin{array}[]{rcl}(B,b)((A,1)(B,b))^{*}&=&(B,b)(A\cup B,b)^{*}\\ &=&(B,b)(b^{-1}A\cup b^{-1}B,1)\\ &=&(B\cup b(b^{-1}A)\cup b(b^{-1}B),b)\\ &=&(B\cup b(b^{-1}A),b),\end{array}$
where the last step follows from the fact that for all $y\in M$ and $X\subseteq M$ we have that $y(y^{-1}X)\subseteq X$. From Remark 3.2, $M$ is a group if and only if this inclusion may always be replaced by $y(y^{-1}X)=X$. Thus if $M$ is a group then $b(b^{-1}A)=A$ and $(A,1)(B,b)=(B,b)((A,1)(B,b))^{*}$. Moreover, we have already seen (in Lemma 3.3) that if $M$ is a group then $\mathcal{S}(M)$ is a proper inverse monoid. Conversely, if $M$ is not a group and we choose $y\in M$ and $X\subseteq M$ with $y(y^{-1}X)\subsetneq X$, then taking $A=X$, $B=\emptyset$ and $b=y$ in the above we see that the right ample identity fails. ∎
Remark 3.6.
We note that Lemma 3.3 together with Proposition 3.4 yields that the expansion $\mathcal{S}$ maps groups to proper inverse monoids, unipotent monoids to proper left restriction monoids, and unipotent right cancellative monoids to left ample monoids. Similarly, by Proposition 3.5 we have that $\mathcal{S}$ maps right abundant monoids to right abundant monoids and left cancellative monoids to right adequate monoids.
4. Finitary conditions
In this section we ask: which monoids $M$ have the property that $\mathcal{S}(M)$ satisfies one of our suite of finitary conditions? In all cases, a necessary (but not sufficient) condition is that $M$ itself satisfies the same finitary property.
4.1. Conditions (R) and (L)
For condition (R) and condition (L), the characterisation turns out to be very simple:
Proposition 4.1.
Let $M$ be a monoid. The following are equivalent:
  1. (1)
    $\mathcal{S}(M)$ satisfies condition (R);
  2. (2)
    $\mathcal{S}(M)$ satisfies condition (L);
  3. (3)
    $M$ is finite.
Proof.
If $M$ is finite then $\mathcal{S}(M)$ is also finite and hence (R) and (L) hold.
To show that (1) implies (3), we prove the contrapositive. Suppose $M$ is infinite and consider the subact $R:=\mathbf{R}((M,1),(M,1))$. By definition
$R$ $=$ $\{((D,d),(E,e)):(M,1)(D,d)=(M,1)(E,e)\}=\{((D,d),(E,e)):M\cup D=M\cup E,d=e\}$
$=$ $\{((D,d),(E,d)):D,E\subseteq M\}.$
We show that $R$ is not finitely generated as a subact of $\mathcal{S}(M)\times\mathcal{S}(M)$. Suppose for contradiction that $X\subseteq R$ is a finite generating set. Let $b,c\in M$ be distinct elements such that $\{b\}\neq P$ and $\{c\}\neq Q$ for all $((P,p),(Q,p))\in X$. Since $((\{b\},1),(\{c\},1))\in R$, there must exist $((P,p),(Q,q))\in X$ and $(T,t)\in\mathcal{S}(M)$ such that
$\begin{array}[]{rcccl}(\{b\},1)&=&(P,p)(T,t)&=&(P\cup pT,pt)\\ (\{c\},1)&=&(Q,p)(T,t)&=&(Q\cup pT,pt).\end{array}$
Since by assumption $\{b\}\neq P$ and $\{c\}\neq Q$, it follows that $\{b\}=pT=\{c\}$, giving the desired contradiction.
Similarly, to show that (2) implies (3), we prove the contrapositive. Suppose $M$ is infinite with cardinality $\kappa$ and consider the subact $L:=\mathbf{L}((M,1),(M,1))$. By definition
$L$ $=$ $\{((D,d),(E,e)):(D,d)(M,1)=(E,e)(M,1)\}=\{((D,d),(E,e)):D\cup dM=E\cup eM,d=e\}$
$=$ $\{((D,d),(E,d)):D\cup dM=E\cup dM\}.$
We show that $L$ is not finitely generated as a subact of $\mathcal{S}(M)\times\mathcal{S}(M)$. Suppose for contradiction that $X\subseteq L$ is a finite generating set. Let
$Y=\{tP:t\in M,\mbox{ and there exists }((P,p),(Q,p))\in X\mbox{ or }((Q,p),(P,p))\in X\}.$
Since $X$ is finite, $|Y|\leq\kappa<2^{\kappa}=|\mathcal{P}(M)|$. So there exists $B\in\mathcal{P}(M)$ such that $B\not\in Y$. Since $((B,1),(M\setminus B,1))\in L$, there must exist $((P,p),(Q,q))\in X$ and $(T,t)\in\mathcal{S}(M)$ such that
$\begin{array}[]{rcccl}(B,1)&=&(T,t)(P,p)&=&(T\cup tP,tp)\\ (M\setminus B,1)&=&(T,t)(Q,p)&=&(T\cup tQ,tp).\end{array}$
Since $B\cap(M\setminus B)=\emptyset$, it follows that $T=\emptyset$ and hence $B=tP\in Y$, giving the desired contradiction. ∎
Remark 4.2.
Note that in Proposition 4.1, to ensure that $M$ is finite, it is enough to insist that either $\mathbf{R}((M,1),(M,1))$ or $\mathbf{L}((M,1),(M,1))$ is finitely generated as a right (respectively, left) act. This certainly implies that $\mathbf{r}((M,1))$ (respectively $\mathbf{l}((M,1))$) is finitely generated as a right (respectively, left) congruence. We will return to this point at the end of the paper.
4.2. The ideal Howson properties
Since $M$ is a retract of $\mathcal{S}(M)$, it follows immediately from [4, Corollary 3.5] and its left-right dual that if $\mathcal{S}(M)$ is right (respectively, left) ideal Howson, then $M$ is right (respectively, left) ideal Howson. On the other hand, if $M$ is left or right ideal Howson it need not be the case that $\mathcal{S}(M)$ is left or right ideal Howson, as we shall demonstrate below. We give a stronger property which will turn out to exactly characterise the monoids $M$ for which $\mathcal{S}(M)$ is right ideal Howson.
Definition 4.3.
(Strongly right ideal Howson) A monoid $M$ is strongly right ideal Howson if for all $a,b\in M$ with $aM\cap bM\neq\emptyset$ there exists $n\geq 1$ and $u_{1},\ldots,u_{n}\in aM\cap bM$ such that
  1. (1)
    $aM\cap bM=\bigcup_{1\leq i\leq n}u_{i}M$;
  2. (2)
    for each $i=1,\ldots,n$, the set $L_{i}=(aM\cap bM)\setminus u_{i}M$ is finite.
If $M$ is strongly right ideal Howson and $aM\cap bM\neq\emptyset$ we say that elements $u_{1},\ldots,u_{n}$ satisfying conditions (1) and (2) above are a strong finite generating set for $aM\cap bM$.
Remark 4.4.
(Principally right ideal Howson) Notice that if $M$ is strongly right ideal Howson and moreover for all $a,b\in M$ with $aM\cap bM\neq\emptyset$ we have a strong finite generating set of cardinality $n=1$, then the first condition simply says that any non-empty intersection of principal ideals is again principal, whilst the second condition is vacuous (the difference is empty). In this case we shall say that $M$ is principally right ideal Howson.
From the definitions it is clear that principally right ideal Howson implies strongly right ideal Howson which in turn implies right ideal Howson. The converse implications do not hold:
Example 4.5.
The construction in [3, Theorem 3.10] may be invoked to give a finite band $B_{2}^{1}$ that is not principally right ideal Howson; since it is finite, it is certainly right ideal Howson. In a similar vein, the semigroup $S=S_{2,0}$ defined in Section 5.1 of [3] with presentation
$S=\langle a,b,w_{1},w_{2},v_{1},v_{2}\mid aw_{1}=bv_{1},aw_{2}=bv_{2}\rangle$
is right ideal Howson and by [3, Lemma 2.2] so too is the monoid $M=S^{1}$; we claim $M$ is not strongly right ideal Howson. To this end, notice that
$a]M\cap[b]M=[aw_{1}]M\cup[aw_{2}]M$
and clearly
$\{[aw_{1}a^{n}]:n\in\mathbb{N}\}$
is an infinite subset of
$([a]M\cap[b]M)\setminus[aw_{2}]M.$
Theorem 4.6.
The semidirect product $\mathcal{S}(M)$ is right ideal Howson if and only if $M$ is strongly right ideal Howson. Moreover, $\mathcal{S}(M)$ is principally right ideal Howson if and only if $M$ is principally right ideal Howson.
Proof.
In what follows we shall suppress the dependence on $M$ and write simply $S:=\mathcal{S}(M)$. Suppose first that $M$ is strongly right ideal Howson. Let $(A,a),(B,b)\in S$ be such that $(A,a)S\cap(B,b)S\neq\emptyset$. Then there exists $(X,x),(Y,y)\in S$ with $(A,a)(X,x)=(B,b)(Y,y)$, that is, $A\cup aX=B\cup bY$ and $ax=by$. In particular, we have that $B\setminus A\subseteq aM$, $A\setminus B\subseteq bM$ and $aM\cap bM\neq\emptyset$. Let $P,Q\subseteq M$ be such that $B\setminus A=aP$ and $A\setminus B=bQ$. Our assumption is that $M$ is strongly right ideal Howson, so there exists a strong finite generating set $u_{1},\ldots,u_{n}$ for $aM\cap bM$. Since each $u_{i}\in aM\cap bM$ we may write $u_{i}=ap_{i}=bq_{i}$ for some $p_{i},q_{i}\in M$. Let $L_{i}$ be defined as in Definition 4.3. We now claim that the finite set
$K:=\{(A\cup B\cup C,u_{i}):C\subseteq L_{i},1\leq i\leq n\}$
is a generating set for $I:=(A,a)S\cap(B,b)S$.
First note that if $C\subseteq L_{i}$ for some $i$, then in particular $C\subseteq aM\cap bM$ and so we may write $C=aC_{a}=bC_{b}$ for some subsets $C_{a},C_{b}$ of $M$, from which it is clear that
$(A,a)(P\cup C_{a},p_{i})$ $=$ $(A\cup aP\cup aC_{a},ap_{i})=(A\cup(B\setminus A)\cup C,u_{i})=(A\cup B\cup C,u_{i}),\mbox{and}$
$(B,b)(Q\cup C_{b},q_{i})$ $=$ $(B\cup bQ\cup bC_{b},bq_{i})=(B\cup(A\setminus B)\cup C,u_{i})=(A\cup B\cup C,u_{i}),$
demonstrating that $K\subseteq I$.
Now suppose that $(D,d)\in I$. Then there exist $(U,u),(V,v)\in S$ such that $(D,d)=(A,a)(U,u)=(B,b)(V,v)$. That is, $D=A\cup aU=B\cup bV$ and $d=au=bv$. Since $d\in aM\cap bM$ we also have $d=u_{i}m$ for some $i$ with $1\leq i\leq n$ and some $m\in M$. Notice that if $c\in D\setminus(A\cup B)$ then $c\in aU\cap bV\subseteq aM\cap bM=\bigcup_{1\leq j\leq n}u_{j}M$. Now let $C=\{c\in D\setminus(A\cup B):c\not\in u_{i}M\}\subseteq L_{i}$, and (noting that $(D\setminus(A\cup B))\setminus C\subseteq u_{i}M$) let $D^{\prime}$ be such that $u_{i}D^{\prime}=(D\setminus(A\cup B))\setminus C$. Then
$(D,d)=(A\cup B\cup C,u_{i})(D^{\prime},m),$
Thus $(D,d)$ is in the right ideal generated by $K$ and it follows that $I=KS$ is finitely generated.
In the case where $M$ is principally right ideal Howson, it follows immediately from the description of the generating set $K$ given above that a non-empty intersection $(A,a)S\cap(B,b)S$ is equal to the principal ideal generated by $(A\cup B,u)$ where $aM\cap bM=uM$, and so $S$ is principally right ideal Howson.
Conversely, suppose that $S$ is right ideal Howson. As commented, [4, Corollary 3.5] gives that a retract of a right ideal Howson monoid is right ideal Howson; the proof of this result also shows that a retract of a principally right ideal Howson monoid is principally right ideal Howson. Thus $M$ is right ideal Howson and principally right ideal Howson if $S$ is. Suppose for contradiction that $M$ is not strongly right ideal Howson. Then there exist $a,b\in M$ such that $aM\cap bM\neq\emptyset$ is finitely generated, but has no strong finite generating set. Fix a generating set $u_{1},\ldots,u_{n}$ for $aM\cap bM$ of $\mathcal{R}$-incomparable elements. For our given generating set, let us also assume that $L_{i}=(aM\cap bM)\setminus u_{i}M$ is infinite (as must be the case for some $i$, since no finite generating set is strong). Let us write $u_{j}=ap_{j}=bq_{j}$ for each $j$. For each $w\in L_{i}$ we may also write $w=u_{j}v$ for some $j\neq i$ and $v\in M$, from which it is clear that
$(\{w\},u_{i})=(\emptyset,a)(\{p_{j}v\},p_{i})=(\emptyset,b)(\{q_{j}v\},q_{i})\in(\emptyset,a)S\cap(\emptyset,b)S.$
Now let $N$ be a generating set for $J:=(\emptyset,a)S\cap(\emptyset,b)S$. With any $w$ as above, $(\{w\},u_{i})\in J$, so there exists $(C,c)\in N$ and $(D,d)\in S$ such that
$(\{w\},u_{i})=(C,c)(D,d)=(C\cup cD,cd).$
Since $(C,c)\in J$ we have in particular that $c\in aM\cap bM=\bigcup_{1\leq j\leq n}u_{j}M$. Moreover, since $u_{i}=cd$ and the generators are $\mathcal{R}$-incomparable we must have $c\in u_{i}M$, and hence $cD\subseteq u_{i}M$. But now, $w\not\in u_{i}M$ and $\{w\}=C\cup cD$ forces $C=\{w\}$. Thus for every $w\in L_{i}$, any generating set of $J$ must contain an element of the form $(\{w\},c)$ where $c\in aM\cap bM$, and since $L_{i}$ is infinite, this demonstrates that each generating set of $J$ is infinite, giving the desired contradiction. ∎
Having characterised the monoids $M$ for which $\mathcal{S}(M)$ is right ideal Howson, we give below a characterisation of the monoids $M$ for which $\mathcal{S}(M)$ is left ideal Howson. This is not quite as neat as the characterisation of right ideal Howson; one requires a condition concerning the left action of $M$ on arbitrary subsets of $M$.
Definition 4.7.
(Left co-ordinated) We say that $a,b\in M$ has finite left co-ordinate system $C$ with respect to subsets $A,B\subseteq M$ if $C$ is a finite subset of the left subact
$\mathbf{L}(a,b)=\{(p,q):pa=qb\}$
of $M\times M$ with the property that for all $x,y\in M$ with $xa=yb$ there exists $t\in M$ and $(p,q)\in C$ such that $xa=tpa=tqb=yb$ and $tpA\cup tqB\subseteq xA\cup yB$. We say that $M$ is (n-)left co-ordinated if for all elements $a,b$ and all subsets $A,B\subseteq M$ there is a finite left co-ordinate system $C$ (with $|C|\leq n)$ for $a,b$ with respect the sets $A,B$.
Remark 4.8.
  1. (1)
    If condition (L) holds, then we also have $M$ is left co-ordinated: taking $C$ to be a finite generating set for the $M$-act $\mathbf{L}(a,b)$, if $xa=yb$, then $(x,y)=t\cdot(p,q)$ for some $(p,q)\in C$ and $t\in M$ hence giving $xa=tpa=tqb=yb$ and $tpA\cup tqB=xA\cup yB$ for all sets $A,B\subseteq M$.
  2. (2)
    If $M$ is left co-ordinated, then it is in particular left ideal Howson: if $C$ is a finite left co-ordinate system for the pair $a,b$ with respect to (any choice of) sets $A$, $B$ then for any $f=xa=yb\in Ma\cap Mb$ we have $f=tpa=tqb$ for some $t\in M$ and $(p,q)\in C$, thus $D=\{pa:(p,q)\in C\}$ is a finite generating set for $Ma\cap Mb$.
  3. (3)
    It follows from (1) that the property of being left co-ordinated is a finitary condition, since this is true of condition (L).
  4. (4)
    Note that $\mathbf{L}(a,a)$ is a left congruence, namely the left annihilator congruence of $a$, and being finitely generated as a left $M$-act certainly implies being finitely generated as a left congruence. Thus if condition (L) holds (i.e. each left $M$-act $\mathbf{L}(a,b)$ is finitely generated) we have in particular that $M$ is finitely left-equated (i.e. each left annihilator congruence $\mathbf{l}(a)$ is finitely generated).
  5. (5)
    From (1) and (2) we have:
[…]
For right cancellative monoids these three properties coincide.
Proposition 4.9.
Let $M$ be a monoid with the property that for all $a,b,f\in M$ the pair of equations $xa=f$, $yb=f$ has at most one solution $(x,y)\in M\times M$. Then: $M$ is left ideal Howson if and only if $M$ is left co-ordinated if and only if $M$ satisfies (L).
Proof.
From the fifth point of Remark 4.8 it suffices to show that if $M$ is left ideal Howson then $M$ satisfies (L). Indeed, if $X$ is a finite generating set for $Ma\cap Mb$, first note that for each $g\in X$ there exists a unique pair $(p_{g},q_{g})\in M\times M$ such that $g=p_{g}a=q_{g}b$, and so taking $C=\{(p_{g},q_{g}):g\in X\}$ we see that $C\subseteq\mathbf{L}(a,b)$. If $f=xa=yb$ holds, then $f=tg$ for some $t\in M$ and $g\in X$, and so $f=tp_{g}a=tq_{g}b$. Then by uniqueness of solution we have $x=tp_{g}$ and $y=tq_{g}$, demonstrating that $C$ is a finite generating set for $\mathbf{L}(a,b)$. ∎
However, in general, this is not the case, as the next two examples demonstrate.
Example 4.10.
We make use of [12, Example 3.8] to demonstrate that a left ideal Howson monoid need not be left co-ordinated. Consider the four element diamond lattice $\mathbb{D}=\{E,F,D,G\}$ where $E\wedge F=G$ and $E\vee F=D$, and let $S$ be the lattice obtained as the direct product
$S=\mathbb{N}\times\mathbb{D}=\{E_{i},F_{i},D_{i},G_{i}:i\geq 0\},$
regarded as a semilattice with respect to $\wedge$ and then let $M=S^{1}$ be the monoid obtained by adjoining an identity to $S$. Since $M$ is a semilattice, it is in particular left ideal Howson. We show that the pair $a=E_{0}$, $b=F_{0}$ does not have a finite left co-ordinate system with respect to the sets $A=B=\{1\}$. First note that
$\mathbf{L}(a,b)=\{(F_{i},E_{j}),(F_{i},G_{j}),(G_{i},E_{j}),(G_{i},G_{j}):i,j\geq 0\}.$
Let $C$ be a finite subset of $\mathbf{L}(a,b)$ and let $m,n\in\mathbb{N}$ be maximal such that $(F_{m},Y)\in C$ for some $Y\in M$ and $(X,E_{n})\in C$ for some $X\in M$. If $C$ were a left co-ordinate system for $a,b$ with respect to $A$ and $B$, then, in particular, for each $i>m$ and $j>y$ there must exist $(p,q)\in C$ and $t\in M$ such that $F_{i}E_{0}=tpE_{0}=tqF_{0}=E_{j}F_{0}$ and $\{tp,tq\}\subseteq\{F_{i},E_{j}\}$. For $tp=F_{i}$, we require $p\geq F_{i}>F_{m}$, and contradicting the maximality of $m$. For $tp=E_{j}$ we require $p\geq E_{j}$, contradicting $(p,q)\in\mathbf{L}(a,b)$.
Example 4.11.
Let $M=\mathbb{Z}\cup\{\bot,\top\}$ ordered by $\bot<x<\top$ for all $x\in\mathbb{Z}$. Clearly $M$ is a monoid semilattice with operation $ab={\rm min}(a,b)$, and identity element $\top$. We show that $M$ is left co-ordinated but does not satisfy $(L)$. Let $a,b\in M$, and $A,B\subseteq M$ and set
$C=\begin{cases}\{(\top,\top)\}&\mbox{ if }a=b\\ \{(\top,a),(a,a)\}&\mbox{ if }a<b\\ \{(b,\top),(b,b)\}&\mbox{ if }b<a.\end{cases}$
We claim that $C$ is a finite left co-ordinate system for $a,b$ with respect to $A,B$. To see this, we consider each of the three cases above ($a=b$, $a<b$ and $b<a$) in turn.
(1) If $a=b$ we have
$C=\{(\top,\top)\}\subseteq\mathbf{L}(a,a)=\{(x,y):xa=ya\}=\{(x,y):x,y\geq a\}\cup\{(x,x):x<a\}.$
We aim to show that for all $(x,y)\in\mathbf{L}(a,a)$ there exists $t\in M$ such that $xa=ta=ya$ and $tA\cup tB\subseteq xA\cup yB$. If $A=\emptyset$ we can take $t=y$; if $B=\emptyset$ we can take $t=x$; and if $x=y$ we can take $t=x$. Thus we may assume without loss of generality that $A$ and $B$ are non-empty and $a\leq x<y$ (a symmetric argument will hold for $y<x$). We subdivide into further cases. First, if all $\alpha\in A$ satisfy $\alpha<a$, then taking $t=y>x\geq a$ yields $tA\cup tB=yA\cup yB=A\cup yB=xA\cup yB$, since both $x$ and $y$ are strictly greater than all elements of $A$. Second, if all $\alpha\in A$ satisfy $\alpha<x$ and there exists $\alpha^{\prime}\in A$ with $a\leq\alpha^{\prime}<x$, then taking $t=\alpha^{\prime}>x\geq a$ yields
$tA\cup tB$ $=$ $\alpha^{\prime}A\cup\alpha^{\prime}B=\{\alpha^{\prime}\alpha:\alpha\in A\}\cup\{\alpha^{\prime}\beta:\beta\in B\}$
$=$ $\{\alpha^{\prime}\}\cup\{\gamma\in A:\gamma<\alpha^{\prime}\}\cup\{\delta\in B:\delta<\alpha^{\prime}\}$
$\subseteq$ $\{\alpha^{\prime}\}\cup\{\gamma\in A:\gamma<x\}\cup\{\delta\in B:\delta<y\}=xA\cup yB$
where the final line follows from the fact that $x$ and $y$ are strictly greater than $\alpha^{\prime}\in A$. Finally, if there exists $\alpha^{\prime\prime}\in A$ with $\alpha^{\prime\prime}\geq x$, then taking $t=x$ gives
$tA\cup tB$ $=$ $xA\cup xB=\{x\alpha:\alpha\in A\}\cup\{x\beta:\beta\in B\}$
$=$ $\{x\}\cup\{\gamma\in A:\gamma<x\}\cup\{\delta\in B:\delta<x\}$
$\subseteq$ $\{x\}\cup\{\gamma\in A:\gamma<x\}\cup\{\delta\in B:\delta<y\}=xA\cup yB$
where the final line follows from the fact that $x<y$.
(2) If $a<b$ we have
$C=\{(\top,a),(a,a)\}\subseteq\mathbf{L}(a,b)$ $=$ $\{(x,y):xa=yb\}$
$=$ $\{(x,a):x\geq a\}\cup\{(x,x):x<a\}$
$=$ $\{x(\top,a):x\geq a\}\cup\{x(a,a):x<a\},$
showing that $\mathbf{L}(a,b)$ is finitely generated by $C$, and hence (by Remark 4.8) $C$ is a finite left co-ordinate set for $a,b$ with respect to $A,B$.
(3) Dually, if $b<a$ we have
$C=\{(b,\top),(b,b)\}\subseteq\mathbf{L}(a,b)$ $=$ $\{(x,y):xa=yb\}$
$=$ $\{(b,y):y\geq b\}\cup\{(y,y):y<b\}$
$=$ $\{y(b,\top):y\geq b\}\cup\{y(b,b):y<b\},$
showing that $\mathbf{L}(a,b)$ is finitely generated by $C$, and hence (by Remark 4.8) $C$ is a finite left co-ordinate set for $a,b$ with respect to $A,B$.
Thus $M$ is finitely left co-ordinated. Finally, to see that $M$ does not satisfy (L), suppose for contradiction that $D$ is a finite generating set for the left $M$-act
$\mathbf{L}(\bot,\bot)=\{(x,y):x\bot=y\bot\}=M\times M.$
Since, in particular $(\top,m)\in\mathbf{L}(\bot,\bot)$ for all $m\in M$, taking $m$ so that it does not occur as a second component of any pair in $D$ gives a contradiction, since we require that $(\top,m)=t(p,q)$ for some $t\in M$ and $(p,q)\in D$, but then $\top=tp$ implies $t=p=\top$, and then $q=m$.
Theorem 4.12.
The semidirect product $\mathcal{S}(M)$ is left ideal Howson if and only if $M$ is left co-ordinated. Further, $\mathcal{S}(M)$ is principally left ideal Howson if and only if $M$ is left $1$-co-ordinated.
Proof.
We again suppress the dependence on $M$ and write simply $S:=\mathcal{S}(M)$. Suppose first $M$ is left co-ordinated. Let $J:=S(A,a)\cap S(B,b)\neq\emptyset$, where $(A,a),(B,b)\in S$. We shall show that $J$ is finitely generated. For each $(F,f)\in J$ there exist $(X,x),(Y,y)\in S$ such that
$\begin{array}[]{rcccl}(F,f)&=&(X,x)(A,a)&=&(X\cup xA,xa)\\ &=&(Y,y)(B,b)&=&(Y\cup yB,yb).\end{array}$
Thus $f=xa=yb$ and $xA\cup yB\subseteq F$ holds. Since $M$ is left co-ordinated, $a,b$ has a finite left co-ordinate system $C$ with respect to the sets $A,B$, and so by assumption we have that there exists $t\in M$ and $(p,q)\in C$ such that $f=tpa=tqb$ and $tpA\cup tqB\subseteq xA\cup yB$. Since $(p,q)\in C\subseteq\mathbf{L}(a,b)$, we have $pa=qb$. Setting $u=pa=qb$ we see that
$(F,f)=(F,t)(pA\cup qB,u).$
Noting that
$(pA\cup qB,u)=(qB,p)(A,a)=(pA,q)(B,b)\in J,$
it then follows that $D:=\{(pA\cup qB,u):(p,q)\in C,u=pa=qb\}$ is a finite generating set for $J$ with $1\leq|D|\leq|C|$. Thus if $M$ is left ($1$-)co-ordinated, then $\mathcal{S}(M)$ is (principally) left ideal Howson.
Conversely, suppose that $S$ is left ideal Howson. Let $a,b\in M$ with $Ma\cap Mb\neq\emptyset$ and let $A,B\subseteq M$. We must find a left co-ordinate system of $Ma\cap Mb$ with respect to the sets $A,B$. Let $c,d\in M$ with $ca=db$ and consider $(A,a),(B,b)\in S$. Notice that
$(dB,c)(A,a)=(dB\cup cA,ca)=(cA\cup dB,db)=(cA,d)(B,b),$
so $J:=S(A,a)\cap S(B,b)\neq\emptyset$. Let $\{(U_{i},u_{i}):1\leq i\leq n\}$ be a finite generating set for $J$. Then for each $i$ we must have $u_{i}=p_{i}a=q_{i}b$ for some $p_{i},q_{i}\in M$, and $p_{i}A\cup q_{i}B\subseteq U_{i}$. Now suppose $f=xa=yb$ for some $x,y\in M$. Then
$(xA\cup yB,f)=(xA\cup yB,x)(A,a)=(xA\cup yB,y)(B,b)\in J,$
and so by definition there exists $1\leq i\leq n$ and $(T,t)\in S$ such that
$(xA\cup yB,f)=(T,t)(U_{i},u_{i})=(T\cup tU_{i},tu_{i}).$
Thus, $f=tu_{i}=tp_{i}a=tq_{i}b$ and $tp_{i}A\cup tq_{i}B\subseteq tU_{i}\subseteq xA\cup yB$ so that $C=\{(p_{i},q_{i}):1\leq i\leq n\}$ is a left co-ordinate system for $a,b$ with respect to the set $A,B$. If $\mathcal{S}(M)$ is (principally) left ideal Howson we see that $M$ is left (1-)co-ordinated. ∎
For right cancellative monoids, we obtain the following simplified characterisation.
Corollary 4.13.
Let $M$ be right cancellative. The semidirect product $\mathcal{S}(M)$ is left ideal Howson if and only if $M$ is left ideal Howson.
Proof.
This follows immediately from Proposition 4.9 and 4.12. ∎
To add to the list of properties in Theorem 2.10 we now have:
Proposition 4.14.
If $M$ is left (1-)co-ordinated, and $T$ is a retract of $M$, then $T$ is also left (1-)co-ordinated.
Proof.
Let $\varphi:M\rightarrow T$ be a retraction, and let $a,b\in T\subseteq M$, and $A,B\subseteq T\subseteq M$. Since $M$ is left co-ordinated, there exists a finite left co-ordinate system $C\subseteq M\times M$ for $a,b$ with respect to $A,B$ considered within $M$. In particular, for each $(x,y)\in T\times T$ with $xa=yb$, there exists $(p,q)\in C$ and $t\in M$ such that $xa=tpa=tqb=yb$ and $tpA\cup tqB\subseteq xA\cup yB$. Applying the morphism $\varphi$ (recalling that this fixes all elements of $T$, and that $a,b,x,y,A,B$ all lie in $T$) yields that $xa=(t\varphi)(p\varphi)a=(t\varphi)(q\varphi)b=yb$ and $(t\varphi)(p\varphi)A\cup(t\varphi)(q\varphi)B\subseteq xA\cup yB$. Further, $pa=qb$ gives that $(p\varphi)a=(q\varphi)b$ so that $C^{\prime}=\{(p\varphi,q\varphi):(p,q)\in C\}$ is a (finite) left co-ordinate system with respect to $A,B$ considered within $T$. Clearly, if $|C|=1$, then $|C^{\prime}|=1$ also. ∎
4.3. The finitely equated conditions
Since $M$ is a retract of $\mathcal{S}(M)$ it follows from [4, Corollary 4.12] that a necessary condition for $\mathcal{S}(M)$ to be finitely right (respectively, left) equated is that $M$ is finitely right (respectively, left) equated, however, as we shall see below, the converse does not hold (in either case). We look first at the case of finitely right equated monoids.
Remark 4.15.
Suppose that $M$ is finitely right equated. Given a finite generating set $X$ for $\mathbf{r}(a)$, setting
$Y:=\{((\emptyset,1),(a^{-1}A,1)),((a^{-1}A,1),(\emptyset,1)),((a^{-1}A,p),(a^{-1}A,q)):(p,q)\in X\},$
is sufficient to generate those elements $((U,u),(V,v))$ of $\mathbf{r}((A,a))$ satisfying $U,V\subseteq a^{-1}A$. Indeed, if $u=p_{1}t_{1},q_{1}t_{1}=p_{2}t_{2},\cdots q_{n}t_{n}=v$ is an $X$-sequence for $(u,v)$, then a $Y$-sequence for $((U,u),(V,v))$ is:
$(U,u)=(\emptyset,1)(U,u),\,\,(a^{-1}A,1)(U,u)$ $=$ $(a^{-1}A,p_{1})(\emptyset,t_{1}),$
$(a^{-1}A,q_{1})(\emptyset,t_{1})$ $=$ $(a^{-1}A,p_{2})(\emptyset,t_{2})$
$\vdots$
$(a^{-1}A,q_{n})(\emptyset,t_{n})$ $=$ $(a^{-1}A,1)(V,v),\,\,(\emptyset,1)(V,v)=(V,v).$
In particular, this demonstrates that each right annihilator of the form $\mathbf{r}((M,a))$ is finitely generated. For an arbitrary subset $A\subseteq M$, the challenge then is finite generation of pairs $((U,u),(V,v))$ where $aU\setminus A=aV\setminus A\neq\emptyset$. We show that this is not possible in complete generality, and give a sufficient condition under which the result holds.
The following lemma will be used show that there exist finitely right equated monoids $M$ such that $\mathcal{S}(M)$ is not finitely right equated.
Lemma 4.16.
Let $M$ be an infinite monoid and suppose that there exists $x\in M$ with the following properties:
  • if $p$ is right invertible in $M$ and $xp=xq$ for some $q\in M$, then $p=q$;
  • for any finite subset $A$ of $M$, there exists $u,v\in M\setminus A$ such that $xu=xv$, but $u\neq v$.
Then $\mathcal{S}(M)$ is not finitely right equated.
Proof.
We show that $\mathbf{r}((\emptyset,x))$ is not finitely generated. Suppose for contradiction that $Y$ is a finite symmetric generating set for $\mathbf{r}((\emptyset,x))$. Let $A$ be the union of all sets $P$ such that $|P|=1$ and there exists $((P,p),(Q,q))\in Y$ for some $p,q\in M$ and some $Q\subseteq M$. Since $Y$ is finite, so too is $A$. Let $u,v\in M\setminus A$ with $xu=xv$ and $u\neq v$ and consider the elements $(\{u\},1),(\{v\},1)\in\mathcal{S}(M)$. We have
$(\emptyset,x)(\{u\},1)=(\{xu\},x)=(\{xv\},x)=(\emptyset,x)(\{u\},1).$
Thus, by assumption, there must be a finite $Y$-sequence from $(\{u\},1)$ to $(\{v\},1)$. Suppose this sequence is:
$(\{u\},1)=(P_{1},p_{1})(T_{1},t_{1}),(Q_{1},q_{1})(T_{1},t_{1})=(P_{2},p_{2})(T_{2},t_{2}),\ldots,(Q_{n},q_{n})(T_{n},t_{n})=(\{v\},1),$
where for $i=1,\ldots,n$ we have $((P_{i},p_{i}),(Q_{i},q_{i}))\in Y$ and $(T_{i},t_{i})\in\mathcal{S}(M)$. For all $i$ we have
$xP_{i}=xQ_{i}\mbox{ and }xp_{i}=xq_{i}.$
Looking at the second component of the $Y$-sequence, we also have the sequence:
$1=p_{1}t_{1},q_{1}t_{1}=p_{2}t_{2},\ldots,q_{n}t_{n}=1.$
Since $p_{1}$ is right invertible and $xp_{1}=xq_{1}$, we deduce that $q_{1}=p_{1}$ and $1=p_{2}t_{2}$. By induction we then find that for $1\leq i\leq n$ each $p_{i}$ is right invertible, $p_{i}=q_{i}$ and $1=p_{i}t_{i}$. We now show by induction that for $1\leq i\leq n$ we have $P_{i}=Q_{i}=\emptyset$ and $p_{i}T_{i}=\{u\}$. Looking at the first component of the first equation in our sequence, we have that $\{u\}=P_{1}\cup p_{1}T_{1}$, which forces $|P_{1}|\leq 1$. Since $u\not\in A$ we deduce that $P_{1}=\emptyset$ and $p_{1}T_{1}=\{u\}$. Since $xP_{i}=xQ_{i}$, this also forces $Q_{i}=\emptyset$ and the base case holds. Suppose then that $P_{i}=Q_{i}=\emptyset$ and $p_{i}T_{i}=\{u\}$ for some $i<n$. Then the equation
$(Q_{i},q_{i})(T_{i},t_{i})=(P_{i+1},p_{i+1})(T_{i+1},t_{i+1})$
yields
$P_{i+1}\cup p_{i+1}T_{i+1}=Q_{i}\cup q_{i}T_{i}=\emptyset\cup p_{i}T_{i}=\{u\}.$
Thus $|P_{i+1}|\leq 1$ and since $u\not\in A,$ we find that $P_{i+1}=\emptyset$ (which in turn forces $Q_{i+1}=\emptyset$) and $p_{i+1}T_{i+1}=\{u\}$. But then we have the contradiction
$\{v\}=Q_{n}\cup q_{n}T_{n}=\emptyset\cup p_{n}T_{n}=\{u\}.$
Example 4.17.
We now use of a construction of John Fountain, (see [10]), adapted to our needs, to demonstrate that there is a finitely right equated monoid $M$ satisfying the conditions of the previous lemma. Let $G$ be a finitely generated infinite group and let $x,x^{2},x^{3}$ be distinct symbols not contained in $G$. Consider the set of formal symbols $M=G\cup xG\cup x^{2}G\cup\{x^{3},x^{4}\}$, with product determined as follows (where we write $x^{0}=1$):
[…]
First, we show that $M$ satisfies the conditions of Lemma 4.16 with respect to the element $x$. Indeed, if $p$ is right invertible, then $p\in G$ and $xp=xq$ implies that $xq\in xG$ and moreover $p=q$, giving that the first condition holds. For the second condition, since $G$ is infinite, for any finite subset $A$ there exist distinct elements $u^{\prime},v^{\prime}\in G$ such that $u=x^{2}u^{\prime},v=x^{2}v^{\prime}$ are distinct elements in $x^{2}G\setminus A$ and $xu=x^{3}=xv$.
It remains to show that $M$ is finitely right equated or, in other words, to show that each of the right annihilators of its elements is finitely generated. We denote the identity relation in $M$ by $\iota_{M}$ and the universal relation in $G$ by $\omega_{G}$. The fact that $G$ is finitely generated (as a group) by a (finite) subset $H$, where we can assume $H=H^{-1}$, gives that $\omega_{G}$ is finitely generated as a right congruence by $H\times H$. Fix $g\in G$. Then $\mathbf{r}(g)=\iota_{M}$ is clearly finitely generated by $\emptyset$. Next,
$\mathbf{r}(xg)$ $=$ $\{(b,c):b,c\in M,(xg)b=(xg)c\}$
$=$ $\iota_{M}\cup\{(x^{2}u,x^{2}v):u,v\in G\}\cup\{(x^{3},x^{4}),(x^{4},x^{3})\}.$
Since for any $u,v\in G$ there is a finite $H\times H$-sequence in $G$ from $u$ to $v$, it follows that there is a finite $x^{2}H\times x^{2}H$-sequence from $x^{2}u$ to $x^{2}v$ in $M$. Thus $\mathbf{r}(xg)$ is finitely generated by $\{(x^{2}h,x^{2}k),(x^{3},x^{4}):h,k\in H\}.$ Similarly,
$\mathbf{r}(x^{2}g)$ $=$ $\{(b,c):b,c\in M,(x^{2}g)b=(x^{2}g)c\}$
$=$ $\iota_{M}\cup\{(xu,xv):u,v\in G\}\cup\{(b,c):b,c\in x^{2}G\cup\{x^{3},x^{4}\}\}$
is finitely generated by $\{(xh,xk),(x^{2},x^{4}):h,k\in H\}.$ We also have
$\mathbf{r}(x^{3})$ $=$ $\{(b,c):b,c\in M,\,x^{3}b=x^{3}c\}$
$=$ $\{(b,c):b,c\in G\}\cup\{(b,c):b,c\in xG\cup x^{2}G\cup\{x^{3},x^{4}\}\}$
is finitely generated by $\{(h,k),(x,x^{4}):h,k\in H\}.$ Finally,
$\mathbf{r}(x^{4})$ $=$ $\{(b,c):b,c\in M\}$
is finitely generated by $\{(1,x^{4})\}$.
It can be deduced from Proposition 2.11 and Proposition 3.5 that if $M$ is right abundant, then $\mathcal{S}(M)$ is finitely right equated. This also follows from a short direct argument:
Lemma 4.18.
Let $a,e\in M$ with $a\,\mathcal{L}^{*}\,e=e^{2}$. Then for any subset $A\subseteq M$, the right annihilator congruence $\mathbf{r}((A,a))$ is finitely generated.
Proof.
Note that, since $ae=a$ we have $(A,a)(\emptyset,1)=(A,a)=(A,a)(a^{-1}A,e)$. Let $\rho$ denote the right congruence generated by the pair $((\emptyset,1),(a^{-1}A,e))$; we shall demonstrate that $\rho$ is equal to $\mathbf{r}((A,a))$. If $((U,u),(V,v))\in\mathbf{r}((A,a))$, then we have $A\cup aU=A\cup aV$ and $au=av$. Moreover, if $x\in eU$ and $x\not\in a^{-1}A$, then $ax\in aU\setminus A=aV\setminus A\subseteq aV=aeV$, but then $ax=aev^{\prime}$ for some $v^{\prime}\in V$ and since $a\,\mathcal{L}^{*}\,e$ we obtain $ex=ev^{\prime}$. Since $x\in eU$ note that $ex=x$ and so $eU\subseteq eV$, and by a dual argument we deduce that $eU=eV$. It follows that
$(U,u)=(\emptyset,1)(U,u)\,\rho\,(a^{-1}A,e)(U,u)$ $=$ $(a^{-1}A\cup eU,eu)$
$=$ $(a^{-1}A\cup eV,ev)=(a^{-1}A,e)(V,v)\,\rho\,(\emptyset,1)(V,v)=(V,v).$
Corollary 4.19.
Let $M$ be a right abundant monoid. Then $\mathcal{S}(M)$ is weakly right coherent if and only if $M$ is strongly right ideal Howson.
On the other hand, as we have alluded to earlier, there exist left abundant monoids $M$ such that $\mathcal{S}(M)$ is not left abundant. Analogously to Lemma 4.16 we now give a criterion for a monoid $M$ that prevents $\mathcal{S}(M)$ from being finitely left equated. We then exhibit an example of a semilattice (which is certainly left abundant) satisfying this condition (so that $\mathcal{S}(M)$ is not finitely left equated and so certainly not left abundant).
In what follows where we say that $b$ is indecomposable we mean that the only factors of $b$ are $1$ and $b$.
Lemma 4.20.
Let $M$ be a monoid and let $a\in M$ be such that there are infinitely many distinct pairs of elements $(b,c)\in\mathbf{l}(a)$ where $b$ is indecomposable and $c\not\leq_{\mathcal{R}}b$. Then $\mathbf{l}((\{1\},a))$ is not finitely generated.
Proof.
We begin by observing that for $((P,p),(Q,q))\in X$ we have that $(P,p)(\{1\},a)=(Q,q)(\{1\},a)$ and so $P\cup\{p\}=Q\cup\{q\}$ and $pa=qa$. Suppose for contradiction that $\mathbf{l}((\{1\},a))$ is generated by a finite set $X$, where we may assume that $X$ is symmetric. Choose $(b,c)\in\mathbf{l}(a)$ such that $b$ is indecomposable, $c\not\leq_{\mathcal{R}}b$ and for all $((P,p),(Q,q))\in X$ we have $P\not\subseteq\{1,b,c\}$ and $p\neq b$.
Let $U\subseteq\{1,b,c\}$ and suppose that $(U,b)=(T,t)(P,p)$ where $(T,t)\in\mathcal{S}(M)$ and there is a pair $((P,p),(Q,q))\in X$. We claim that $(T,t)(Q,q)=(V,b)$ where $V\subseteq\{1,b,c\}$. From $(U,b)=(T,t)(P,p)$ we have that $U=T\cup tP$ and $b=tp$. Since $b$ is indecomposable and $p\neq b$ we deduce that $p=1$ and $b=t$. Further, $T\cup bP\subseteq\{1,b,c\}$. Let $r\in P$. Then $br=1,b$ or $c$. The only possibility (since $c\not\leq_{\mathcal{R}}b$) is for $br=b$ so that $r\in\{1,b\}$. We now split the argument into two cases.
If $b$ is not idempotent, then $r=1$ and so $P\subseteq\{1\}$. Since $P\cup\{p\}=Q\cup\{q\}$ we have that $\{1\}=Q\cup\{q\}$ so that $Q\subseteq\{1\}$ and $q=1$. Then $(T,t)(Q,q)=(T,b)(Q,1)=(V,b)$ where $V\subseteq\{1,b,c\}$ as required.
If $b$ is idempotent, then $r\in\{1,b\}$ and so $P\subseteq\{1,b\}$. Since $P\cup\{p\}=Q\cup\{q\}$ we have that $Q\subseteq\{1,b\}$ and $q\in\{1,b\}$. Then, using the fact that $b$ is idempotent, $(T,t)(Q,q)=(T,b)(Q,q)=(V,b)$, where $V\subseteq\{1,b,c\}$ as required. This verifies our claim.
Notice now that $((\{b,c\},b),(\{b,c\},c)\in\mathbf{l}((\{1\},a))$ and as $b\neq c$ there must be an $X$-sequence of positive length connecting $(\{b,c\},b)$ to $(\{b,c\},c)$. But this is now shown to be impossible. We deduce that $\mathbf{l}((\{1\},a))$ is not finitely generated. ∎
Example 4.21.
Let $M$ be the monoid with presentation as a monoid semilattice given by
$M:=\langle a,b_{i},c_{i},i\in\mathbb{N}\mid b_{i}a=c_{i}a\rangle.$
Since $M$ is a semilattice, it is certainly abundant and so finitely left equated. By construction, we have infinitely many distinct pairs $(b_{i},c_{i})\in\mathbf{l}(a)$. It is easily seen that $b_{i}$ (indeed also $c_{i}$ and $a$) is irreducible for all $i\in\mathbb{N}$. In particular, $c_{i}$ is not $\mathcal{R}$-below $b_{i}$. Thus by Lemma 4.20 we see that $M$ is not finitely left equated.
We can build on this example to construct further examples of finitely left equated monoids satisfying the conditions of Lemma 4.20. Take $M$ as above and any cancellative monoid $T$ with trivial group of units (for example, a free monoid). By [4, Corollary 4.26] we have that $M\times T$ is finitely left equated. It is then easy to see that $(a,1_{T})$ satisfies the conditions of Lemma 4.20 so that $\mathcal{S}(M\times T)$ is not finitely left equated.
Whilst left abundance of $M$ is not sufficient to guarantee that $\mathcal{S}(M)$ is finitely left equated, we do have the following sufficient condition.
Proposition 4.22.
Let $M$ be a right cancellative monoid. Then $\mathcal{S}(M)$ is finitely left equated. Consequently, $\mathcal{S}(M)$ is weakly left coherent if and only if $M$ is left ideal Howson.
Proof.
Suppose $M$ is right cancellative. For each $\alpha=(A,a)\in\mathcal{S}(M)$, we show that the left congruence $\mathbf{l}(\alpha)=\lambda_{W}$ where $W=\{(A,1),(\emptyset,1))\}$. Certainly $W\subseteq\mathbf{l}(\alpha)$, since:
$(A,1)(A,a)=(A\cup A,a)=(A,a)=(\emptyset\cup A,a)=(\emptyset,1)(A,a).$
Notice also that for any subsets $R,T$ of $A$ we have $(T,1)\,\lambda_{W}\,(R,1)$ since:
$(T,1)=(T,1)(\emptyset,1)\,\lambda_{W}\,(T,1)(A,1)=(A,1)=(R,1)(A,1)\,\lambda_{W}\,(R,1)(\emptyset,1)=(R,1).$
Now for $((U,u),(V,v))\in\mathbf{l}(\alpha)$, by definition, we have
$(U\cup uA,ua)=(U,u)(A,a)=(V,v)(A,a)=(V\cup vA,va),$
and by right cancellativity of $M$ it follows that $u=v$ and $U\cup uA=V\cup uA$. Let $U^{\prime}=U\setminus(U\cap uA)$ and $V^{\prime}=V\setminus(V\cap uA)$ so that $U^{\prime}=V^{\prime}$ and let $R,T\subseteq A$ with $uR=U\cap uA$ and $uT=V\cap uA$. Then
$(U,u)=(U^{\prime},u)(R,1)\,\lambda_{W}\,(U^{\prime},u)(T,1)=(V^{\prime},v)(T,1)=(V,v).$
Thus for $M$ right cancellative we have that $\mathcal{S}(M)$ is weakly left coherent if and only if it is left ideal Howson, which by Proposition 4.12 is equivalent to $M$ being left co-ordinated. Since $M$ is right cancellative Proposition 4.9 gives that this is equivalent to the condition that $M$ is left ideal Howson. ∎
Recalling from [7] that a monoid is left (right) LCM if it is right (left) cancellative and principally left (right) ideal Howson, we have:
Corollary 4.23.
If $M$ is left (right) LCM, then $\mathcal{S}(M)$ is weakly left (right) coherent.
Proof.
The result follows directly from Proposition 4.22 and Corollary 4.19. ∎
The class of left LCM monoids includes: groups, free monoids, the multiplicative monoid of non-zero elements in any LCM domain, and Artin monoids (see [7] for details).
The article focuses exclusively on the expansion $\mathcal{S}(M)$ of a monoid $M$. As indicated in the Introduction, this is the largest of a suite of possible expansions, where we make various restrictions on the pairs $(A,a)\in\mathcal{P}(M)\times M$ in the expansion. These include the restriction that $A$ is finite, resulting in the corresponding expansion $\mathcal{S}^{f}(M)$, and the restriction that $A$ is a finite set containing $\{1,a\}$ as a subset, resulting in the Szendrei expansion Sz(M)(M). In the latter case the expansion is to a monoid subsemigroup of $\mathcal{S}(M)$ with identity $(\{1\},1)$.
Question 4.24.
What conditions must a monoid $M$ satisfy such that $\mathcal{S}^{f}(M)$, Sz(M)(M), or other natural related expansions, have the finitary conditions investigated in this article?
We remark that the techniques we have developed here (which for example make frequent use of the fact that the sets $A$ can be chosen to be empty or infinite) in some cases would not apply to these ‘smaller’ expansions so that new strategies would be required.
For a right congruence $\rho_{W}$ (respectively, a left congruence $\lambda_{W}$) on a monoid $M$, we may call the sequence $c_{1},d_{1},\cdots,c_{n}d_{n}$ appearing in the (left-right dual of the) $W$-sequence in Proposition 2.1 the skeleton of the $W$-sequence. If $M$ is right abundant, then we have shown in Lemma 4.18 that $\mathcal{S}(M)$ is finitely right equated and, moreover, for any fixed $(A,a)\in\mathcal{S}(M)$ we may choose a finite (indeed, singleton) generating set $W$ for $\mathbf{r}((A,a))$ such that any two elements of $\mathbf{r}((A,a))$ are connected by a $W$-sequence of length 2 and with the same skeleton. Likewise, if $M$ is right cancellative, then Proposition 4.22 yields that $\mathcal{S}(M)$ is finitely left equated and, moreover, for any fixed $(A,a)\in\mathcal{S}(M)$ we may choose a finite (indeed, singleton) generating set $W$ for $\mathbf{l}((A,a))$ such that any two elements of $\mathbf{l}((A,a))$ are connected by a $W$-sequence of length 2 and with the same skeleton. On the other hand, if condition (L) holds for $\mathcal{S}(M)$ so that in particular $\mathbf{l}((A,a))$ is finitely generated as a left $\mathcal{S}(M)$-act by a subset $W$, then any two elements of $\mathbf{l}((A,a))$ are connected by a $W$-sequence of length 1. This leads us to the following questions.
Question 4.25.
Let $M$ be a monoid. What conditions must $M$ satisfy such that for any $(A,a)\in\mathcal{S}(M)$ we have that $\mathbf{l}((A,a))$ is finitely generated by a set $W$ such that (a) only one skeleton is required (b) there is a bound on the length of $W$-sequences?